# Addressing Data Heterogeneity in Federated Learning with Adaptive Normalization-Free Feature Recalibration

### v1

@@ Abstract
Federated learning is a decentralized collaborative training paradigm that preserves stakeholders’ data ownership while improving performance and generalization. However, statistical heterogeneity among client datasets poses a fundamental challenge by degrading system performance. To address this issue, we propose Adaptive Normalization-free Feature Recalibration (ANFR), an architecture-level approach that combines weight standardization and channel attention. Weight standardization normalizes the weights of layers instead of activations. This is less susceptible to mismatched client statistics and inconsistent averaging, thereby more robust under heterogeneity. Channel attention produces learnable scaling factors for feature maps, suppressing those that are inconsistent between clients due to heterogeneity. We demonstrate that combining these techniques boosts model performance beyond their individual contributions, by enhancing class selectivity and optimizing channel attention weight distribution. ANFR operates independently of the aggregation method and is effective in both global and personalized federated learning settings, with minimal computational overhead. Furthermore, when training with differential privacy, ANFR achieves an appealing balance between privacy and utility, enabling strong privacy guarantees without sacrificing performance. By integrating weight standardization and channel attention in the backbone model, ANFR offers a novel and versatile approach to the challenge of statistical heterogeneity. We demonstrate through extensive experiments that ANFR consistently outperforms established baselines across various aggregation methods, datasets, and heterogeneity conditions.
@@ 1 Introduction
Federated learning (FL) (McMahan et al., 2017) is a decentralized training paradigm which enables clients to jointly develop a global model without sharing private data. By preserving data privacy and ownership, FL holds significant promise for applications across various domains, including healthcare, finance, and mobile devices. However, a fundamental challenge in FL is statistically heterogeneous, i.e. non-independent and identically distributed (non-IID) client datasets. This heterogeneity can degrade the performance of the global model and slow down convergence (Li et al., 2020b; Hsu et al., 2019). Addressing it is thus critical for the success of FL in real-world scenarios.
Most prior relevant research has focused on developing novel aggregation algorithms that either aim to reduce client drift (Karimireddy et al., 2020) or tailor specific layers of the model to individual clients to accommodate data variations (Zhang et al., 2023). While these algorithmic strategies have advanced the field, they often overlook the impact of the underlying model architecture on performance under heterogeneity. Of particular interest is the role of normalization layers. Batch Normalization (BN) (Ioffe & Szegedy, 2015), although effective in centralized settings, has been shown to negatively affect performance in heterogeneous FL due to mismatched client-specific statistics and inconsistent parameter averaging (Wang et al., 2023; Guerraoui et al., 2024). In response, the use of alternative feature normalization methods like Group Normalization (GN) (Wu & He, 2018) and Layer Normalization (LN) (Ba et al., 2016) has been frequent (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021; Du et al., 2022). While these alternatives do not rely on mini-batch statistics, they offer inconsistent benefits, often resulting in slower convergence and lower performance compared to BN (Chen & Chao, 2021; Tenison et al., 2023; Zhong et al., 2024). Previous works have not designed models specifically tailored to combat heterogeneity, leaving a gap in FL research.
@@
We address this gap in the image domain by proposing Adaptive Normalization-Free Feature Recalibration (ANFR), an architecture-level approach designed to enhance robustness in FL under data heterogeneity. ANFR combines weight standardization (Qiao et al., 2020) with channel attention mechanisms (Hu et al., 2018) to directly tackle the challenges posed by non-IID data. Weight standardization normalizes the weights of convolutional layers instead of activations, eliminating the dependence on mini-batch statistics that are problematic in FL. This makes the model less susceptible to mismatched client statistics and inconsistent averaging during aggregation. Channel attention modules produce learnable scaling factors for feature maps, enabling the model to suppress features that are inconsistent across clients due to heterogeneity and to emphasize those that are consistent and relevant. By integrating channel attention with weight-standardized models, ANFR enhances the model’s ability to focus on shared, informative features across clients. This combination boosts performance beyond the individual contributions of weight standardization and channel attention, enhancing class selectivity and optimizing the distribution of channel attention weights. ANFR operates independently of the aggregation method and is effective in both global and personalized FL settings, with minimal computational overhead. Furthermore, when training with differential privacy, ANFR achieves an appealing balance between privacy and utility, enabling strong privacy guarantees without sacrificing performance.
@@
We validate the effectiveness of ANFR through extensive experiments on a diverse set of datasets, including medical imaging and natural image classification tasks, under various types of data heterogeneity. Our results demonstrate that ANFR consistently outperforms established baselines across different aggregation methods, datasets, and heterogeneity conditions. By focusing on architectural components, our approach is orthogonal to existing algorithmic advancements and represents a crucially underdeveloped area in FL research. The proposed ANFR model offers a robust and flexible solution to the challenge of statistical heterogeneity, contributing to the advancement of federated learning by enhancing performance, stability, and privacy-preserving capabilities.
@@ 2 Related Work
Since McMahan et al. (2017) introduced FL, most research has focused on developing novel aggregation algorithms to address challenges like data heterogeneity. In global FL (GFL), methods such as proximal regularization (Li et al., 2020a) and cross-client variance reduction (Karimireddy et al., 2020) aim to reduce client drift. Techniques like discouraging dimensional collapse through correlation matrix norm regularization (Shi et al., 2023), adopting relaxed adversarial training (Zhu et al., 2023), and performing amplitude normalization in frequency space (Jiang et al., 2022) have also been proposed. Recent promising ideas include constructing global pseudo-data to de-bias local classifiers and features (Guo et al., 2023), introducing concept drift-aware adaptive optimization (Panchal et al., 2023), and using hyperbolic graph manifold regularizers (An et al., 2023). In personalized FL (pFL), personalizing specific layers of the model can mitigate heterogeneity. The simplest approach shares all model parameters except the classification head (Arivazhagan et al., 2019). More advanced methods replace lower layers and mix higher ones (Zhang et al., 2023) or adjust mixing ratios based on convergence rate approximations (Jiang et al., 2024). While these algorithmic approaches have advanced both GFL and pFL, they often overlook the impact of the underlying architecture on performance.
@@
Our work addresses this gap by examining how specific model components can enhance FL performance. This approach is orthogonal to algorithmic advancements, representing a crucially underdeveloped area. Previously, Qu et al. (2022) demonstrated performance gains by switching from a pre-trained convolutional network to a pre-trained vision transformer. Studies by Pieri et al. (2023) and Siomos et al. (2024) evaluated multiple architectures under different aggregation methods, showing that changing the backbone model can improve performance more than altering the aggregation method. However, these works did not design models specifically tailored to combat heterogeneity. In contrast, our method directly addresses data heterogeneity by integrating into the model architectural components that enhance robustness across diverse client distributions.
@@
The normalization layer has been a focal point of component examination because Batch Normalization (BN) (Ioffe & Szegedy, 2015), the most widely used normalization technique, has been shown both theoretically (Li et al., 2021; Wang et al., 2023) and empirically (Hsieh et al., 2020; Du et al., 2022; Guerraoui et al., 2024) to negatively impact performance in heterogeneous FL settings. Mismatched local distributions lead to averaged batch statistics and parameters that fail to accurately represent any source distribution. Two primary approaches address this issue: modifying the aggregation rule for the BN layer or replacing it entirely. Some methods keep BN parameters local to each client (Li et al., 2021; Andreux et al., 2020) or stop sharing them after a certain round (Zhong et al., 2024). Others replace batch-specific statistics with shared running statistics when normalizing batch inputs to match local statistical parameters (Guerraoui et al., 2024) or leverage layer-wise aggregation to also match associated gradients (Wang et al., 2023). These methods still rely on decently sized batches to accurately approximate statistics and are incompatible with differential privacy. To replace BN, Group Normalization (GN) (Wu & He, 2018) has been frequently used (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021) since it does not rely on mini-batch statistics. Du et al. (2022) showed that Layer Normalization (LN) (Ba et al., 2016) performs better than GN in some FL settings. However, studies have shown these alternative normalization techniques offer inconsistent benefits over BN, depending on factors like dataset characteristics and heterogeneity levels (Tenison et al., 2023; Chen & Chao, 2021; Zhong et al., 2024). Moreover, GN requires tuning the number of groups for optimal effectiveness.
We propose circumventing these issues by applying weight standardization (Qiao et al., 2020) to normalize the weights of the model instead of the activations. Inspired by Brock et al. (2021a), who showed that such Normalization-Free (NF) models can train stably and perform on par with those employing BN in centralized learning, we explore this concept in FL. Previously, Zhuang & Lyu (2024) proposed an aggregation method specific to NF models for multi-domain FL with small batch sizes. Similarly, Siomos et al. (2024) showed that NF-ResNets improve upon vanilla ResNets under different initialization schemes and aggregation methods, while Kang et al. (2024) proposed a personalized aggregation scheme that replaces each BN layer with weight normalization (Salimans & Kingma, 2016) followed by a learnable combination of BN and GN.
@@
Our method adaptively recalibrates the resulting feature maps using channel attention modules, such as the Squeeze-and-Excitation block (Hu et al., 2018). By doing so, we enhance the model’s ability to focus on relevant features across clients, effectively addressing data heterogeneity. Zheng et al. (2022) previously explored channel attention for FL in a different context, proposing a personalized FL method where a modified channel attention block is kept personal to each client. Unlike previous methods limited to specific aggregation strategies or settings, our approach can complement any heterogeneity-focused aggregation method, is effective even with large batch sizes, and supports various attention modules. Appendix C summarizes the differences between ANFR and related work. By integrating weight standardization with channel attention, ANFR provides a robust and flexible solution to data heterogeneity in FL, overcoming limitations of activation normalization techniques and complementing various aggregation strategies.
@@ 3 Adaptive Normalization-Free Feature Recalibration
We consider a FL setting with clients, each owning a dataset of image-label pairs and optimizing a local objective , where is a loss function and the model parameters. The global objective is learning a model that minimizes the aggregate loss:
Heterogeneity among can degrade the global model performance and slow convergence (Kairouz et al., 2021). In this study, we are interested in modifying the backbone model to address it. As they are the most widely used family, and they perform better or on par with others (Pieri et al., 2023; Siomos et al., 2024), we focus specifically on convolutional neural networks (CNNs). Let represent a batch of image samples with channels and dimensions . For a convolutional layer with weights and a kernel size of , the outputs are given by:
In typical CNNs, the activations are then normalized:
where are learnable parameters, is an indexing vector and is the set of pixels over which are computed. BN computes statistics along the axes, LN along , and GN along separately for each of groups of channels.
Channel attention (CA) mechanisms, like the Squeeze-and-Excitation (SE) block (Hu et al., 2018), recalibrate feature responses by modeling inter-channel relationships. The channel descriptor is obtained via Global Average Pooling (GAP):
This descriptor is then non-linearly transformed to capture dependencies between channels; in SE blocks this is done via the learnable weights and , where is a dimensionality reduction ratio:
yielding per-channel scaling factors which are applied to the normalized activations .
@@
Comparing (6) and (8), we note several advantages of ANFR. First, and are computed from convolutional weights, not the activations. Since weights are initialized identically and synchronized during FL, these weight-derived statistics are consistent across clients. Moreover, the second term of (8) now captures statistics of the input before convolution, providing an additional calibration point for CA and bypassing the effect of . By applying CA after SWS, we ensure channel descriptors are not distorted by batch-dependent statistics or cross-channel dependencies introduced by activation normalization. This allows CA to adjust channel responses effectively, improving the model’s capacity to learn stable feature representations that are consistent across clients with diverse data distributions. Therefore, the combination of SWS and CA overcomes the drawbacks of traditional normalization methods in federated learning, providing a novel and effective solution for improving model performance in the presence of data variability.
@@
To further substantiate the effectiveness of integrating CA with weight standardization in addressing data heterogeneity, we conduct a mechanistic interpretability analysis comparing the effects of BN and SWS on class selectivity and attention weight variability. Specifically, we examine how well the ANFR model discriminates between classes both before111All networks are pre-trained on ImageNet. and after training on the heterogeneous ‘split-3’ partitioning of CIFAR-10 from Qu et al. (2022). This evaluation helps us understand how our method increases class discriminability under data heterogeneity. To isolate the effect of different components, we compare ANFR (using SWS with CA), to BN-ResNet (using BN), NF-ResNet (using SWS without CA) and SE-ResNet (using CA with BN). We quantify class selectivity using the class selectivity index (CSI) (Morcos et al., 2018), defined for each neuron as , where is the class-conditional activation that elicits the highest response and is the mean activation for all other classes. A right-skewed CSI distribution indicates higher class selectivity, crucial for effective classification under heterogeneous data. For models using CA (SE-ResNet and ANFR), we also examine attention weights distribution following Wang et al. (2020) to understand CA’s contribution to class discrimination.
@@
Figure 2 shows CSI distributions for the last layer before the classifier, where class specificity is maximized in CNNs. Prior to FL training, incorporating CA in SE-ResNet with BN results in a small increase in class selectivity compared to BN-ResNet. When combining CA with SWS in ANFR, the change in class selectivity is negligible compared to NF-ResNet, indicating the minimal impact of CA at this stage. After training on heterogeneous data, we observe a notable shift: BN causes reduced class selectivity (compared to before training), evidenced by left-skewed distributions for BN-ResNet and SE-ResNet. Adding CA increases class selectivity for both normalization methods, but cannot fully mitigate BN’s negative effect due to inconsistently normalized inputs to the CA module (6). The ANFR model, however, shows a significant increase in class selectivity compared to NF-ResNet. This discrepancy manifests only after FL training, and indicates that combining CA and SWS in ANFR enhances the model’s ability to specialize in class discrimination under data heterogeneity.
@@
For models using CA, we use the variability of attention weights across channels and classes as an indicator of adaptation: high variability suggests CA is actively re-weighing features to adapt to different class characteristics. Figure 3 shows that before FL training, both SE and ANFR models display high variability. This shows that CA provides a diverse and informative signal for both activation and weight normalization when heterogeneity is not a factor. However, after FL training, the attention weights of SE-ResNet become saturated, converging towards a value of 1 across all channels and classes. This saturation renders the attention mechanism inactive, turning it into an identity operation. Thus, the SE-ResNet model fails to preserve the discriminative power of CA under heterogeneity. Instead, the ANFR model maintains high variability in CA weights across channels and classes. This sustained variability implies that CA remains active and continues to provide class-discriminative signals when combined with weight standardization.
@@
These mechanistic insights support our design choices. The adverse effects of BN under heterogeneous FL conditions are highlighted by the diminished class selectivity and inactive CA in SE-ResNet, while the ability of ANFR to maintain and improve class selectivity demonstrates that integrating CA with weight standardization effectively counters the impact of data heterogeneity. The enhanced class selectivity in ANFR correlates with improved downstream performance in heterogeneous FL settings, as we will show in Section 4. Additional details and extended CSI and attention weight results from other layers are presented in Appendix D.

### v2

# An Architecture Built for Federated Learning: Addressing Data Heterogeneity through Adaptive Normalization-Free Feature Recalibration

@@ Abstract
Federated learning is a decentralized collaborative training paradigm preserving stakeholders’ data ownership while improving performance and generalization. However, statistical heterogeneity among client datasets degrades system performance. To address this issue, we propose Adaptive Normalization-free Feature Recalibration (ANFR), a model architecture-level approach that combines weight standardization and channel attention to combat heterogeneous data in FL. ANFR leverages weight standardization to avoid mismatched client statistics and inconsistent averaging, ensuring robustness under heterogeneity, and channel attention to produce learnable scaling factors for feature maps, suppressing inconsistencies across clients due to heterogeneity. We demonstrate that combining these techniques boosts model performance beyond their individual contributions, by improving class selectivity and channel attention weight distribution. ANFR works with any aggregation method, supports both global and personalized FL, and adds minimal overhead. Furthermore, when training with differential privacy, ANFR achieves an appealing balance between privacy and utility, enabling strong privacy guarantees without sacrificing performance. By integrating weight standardization and channel attention in the backbone model, ANFR offers a novel and versatile approach to the challenge of statistical heterogeneity. Extensive experiments show ANFR consistently outperforms established baselines across various aggregation methods, datasets, and heterogeneity conditions. Code is provided at https://github.com/siomvas/ANFR.
@@ 1 Introduction
Federated learning (FL) (McMahan et al., 2017) is a decentralized training paradigm enabling clients to jointly develop a model without sharing private data. By preserving data privacy and ownership, FL holds promise for applications in healthcare, finance, and mobile devices. A fundamental challenge in FL is statistically heterogeneous, i.e. non-independent and identically distributed (non-IID) client datasets, as they can degrade the performance of the global model and hinder convergence (Li et al., 2020b; Hsu et al., 2019). Addressing this is critical for FL’s success in real-world scenarios. Most prior research focuses on aggregation methods to compensate for this issue, overlooking how model architecture affects performance under heterogeneity. More specifically, Batch Normalization (BN) (Ioffe & Szegedy, 2015) hinders performance in heterogeneous FL due to mismatched client-specific statistics and inconsistent parameter averaging (Wang et al., 2023; Guerraoui et al., 2024). In response, using other feature normalization methods like Group Normalization (GN) (Wu & He, 2018) and Layer Normalization (LN) (Ba et al., 2016) has been frequent in FL research (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021; Du et al., 2022). These alternatives slow convergence and reduce performance compared to BN (Chen & Chao, 2021; Tenison et al., 2023; Zhong et al., 2024). Previous work on designing models specifically tailored to combat heterogeneity is limited to only Xu et al. (2024), leaving a research gap.
@@
We address this gap in the image domain by proposing Adaptive Normalization-free Feature Recalibration (ANFR), an architecture-level approach designed to enhance robustness in FL under data heterogeneity. ANFR combines weight standardization (Qiao et al., 2020) with channel attention (Hu et al., 2018) to directly tackle the challenges posed by non-IID data. Weight standardization normalizes convolutional layer weights instead of activations, avoiding reliance on mini-batch statistics, which is problematic in FL. This reduces susceptibility to mismatched statistics and inconsistent averaging. Channel attention generates learnable scaling factors for feature maps, suppressing features that are inconsistent across clients due to heterogeneity, and emphasizing consistent ones. By integrating channel attention with weight-standardized models, ANFR enhances the model’s ability to focus on shared, informative features across clients. This synergy boosts performance beyond the individual contributions of these components, enhancing class selectivity, and optimizing channel attention weight distribution. A key advantage of ANFR is its versatility; it can be seamlessly integrated with any aggregation method and is highly effective in both global and personalized FL scenarios, while adding minimal computational overhead. Furthermore, when privacy is a concern, our achieves a noteworthy balance between the competing demands of privacy and utility. This enables the implementation of strong privacy guarantees through differential privacy, without a substantial sacrifice in the model’s overall performance.
@@
We validate the effectiveness of ANFR through extensive experiments on a diverse set of datasets and tasks, including medical imaging and natural image classification, multi-class classification, and cross-device scenarios, under various types of data heterogeneity. The results show that ANFR consistently outperforms established baselines across different aggregation methods, datasets, and heterogeneity conditions. By focusing on architectural components, our approach complements advances in aggregation strategies and addresses a crucial gap in FL research. The proposed model offers a robust and flexible solution to the challenge of statistical heterogeneity, contributing to the advancement of federated learning by improving performance, stability, and privacy-preserving capabilities.
@@ 2 Related Work
Since McMahan et al. (2017) introduced FL, most research has focused on developing aggregation algorithms to address challenges like data heterogeneity. In global FL (GFL), methods such as proximal regularization (Li et al., 2020a) and cross-client variance reduction (Karimireddy et al., 2020) aim to reduce client drift. Techniques like discouraging dimensional collapse through correlation matrix norm regularization (Shi et al., 2023), adopting relaxed adversarial training (Zhu et al., 2023), and performing amplitude normalization in frequency space (Jiang et al., 2022) have also been proposed. Other recent ideas are constructing global pseudo-data to de-bias local classifiers and features (Guo et al., 2023), introducing concept drift-aware adaptive optimization (Panchal et al., 2023), and hyperbolic graph manifold regularizers (An et al., 2023). In personalized FL (pFL), personalizing layers of the model can mitigate heterogeneity. The simplest approach shares all model parameters except the classification head (Arivazhagan et al., 2019). More advanced methods replace lower layers and mix higher ones (Zhang et al., 2023) or adjust mixing ratios based on convergence rate approximations (Jiang et al., 2024). While these algorithmic approaches have advanced both GFL and pFL, they often overlook the impact of the underlying architecture on performance.
@@
We address this gap by exploring how model components can enhance FL performance. This is orthogonal to algorithmic advancements, representing a crucially underdeveloped area. Previously, Qu et al. (2022) found using vision transformers instead of convolutional networks increased performance. Studies by Pieri et al. (2023) and Siomos et al. (2024) evaluated different architectures and aggregation methods, showing that changing the architecture, rather than the aggregation method, can be more beneficial. These works did not design models specifically tailored to combat heterogeneity. More recently, Xu et al. (2024) investigated the impact of CNN micro-design choices, such as activation functions and block structure, in FL, and proposed FedConv, an architectural design based on ResNet informed by these investigations. Our method follows in this line of work, namely integrating architectural components that enhance robustness across diverse client distributions into the model, directly addressing data heterogeneity.
@@
The normalization layer has been a focal point of component examination as Batch Normalization (BN) (Ioffe & Szegedy, 2015) has been shown both theoretically (Li et al., 2021; Wang et al., 2023) and empirically (Hsieh et al., 2020; Du et al., 2022; Guerraoui et al., 2024) to negatively impact performance in heterogeneous FL. Mismatched local distributions lead to averaged batch statistics and parameters that fail to accurately represent any source distribution. The primary approaches addressing this issue are modifying the aggregation rule for the BN layer or replacing it entirely. Some methods keep BN parameters local (Li et al., 2021; Andreux et al., 2020) or stop sharing them after a certain round (Zhong et al., 2024). Others replace batch-specific statistics with shared running statistics when normalizing batch inputs to match local statistical parameters (Guerraoui et al., 2024) or leverage layer-wise aggregation to also match associated gradients (Wang et al., 2023). These methods rely on decently sized batches to accurately approximate statistics and are incompatible with differential privacy. To replace BN, Group Normalization (GN) (Wu & He, 2018) has been frequently used (Hsieh et al., 2020; Reddi et al., 2021; Wang et al., 2021) since it does not rely on mini-batch statistics. However, tuning the number of groups in GN is required to maximize effectiveness and Du et al. (2022) showed that Layer Normalization (LN) (Ba et al., 2016) performs better than GN in some settings. Separate studies have shown both GN and LN offer inconsistent benefits over BN, depending on the characteristics and heterogeneity of the dataset (Tenison et al., 2023; Chen & Chao, 2021; Zhong et al., 2024).
@@
We circumvent these issues by applying weight standardization (Qiao et al., 2020) to normalize the weights of the model instead of the activations. Inspired by Brock et al. (2021a), who showed that such Normalization-Free (NF) models can train stably and perform on par with BN in centralized learning, we explore this concept in FL. Previously, Zhuang & Lyu (2024) proposed an aggregation method specific to NF models for multi-domain FL with small batch sizes. Similarly, Siomos et al. (2024) showed that NF-ResNets improve upon vanilla ResNets under different initialization schemes and aggregation methods, while Kang et al. (2024) proposed a personalized aggregation scheme that replaces each BN layer with weight normalization (Salimans & Kingma, 2016) followed by a learnable combination of BN and GN. Additionally, our method adaptively recalibrates the resulting feature maps using channel attention modules, such as the Squeeze-and-Excitation block (Hu et al., 2018). By doing so, the model can focus more on relevant features across clients, effectively addressing data heterogeneity. Zheng et al. (2022) previously explored channel attention for pFL, proposing a modified channel attention block that is kept personal to each client. Unlike previous methods limited to specific aggregation strategies or settings, our approach can complement any heterogeneity-focused aggregation method, is effective even with large batch sizes, and supports various attention modules. By integrating weight standardization with channel attention, ANFR provides a robust and flexible solution to data heterogeneity in FL, overcoming limitations of activation normalization techniques and complementing aggregation methods. Table 1 presents a tabular comparison of ANFR with related work to make it easier to understand our contributions.
Method | Scenario |
GFL | pfL | ||||||||||
ANFR (Ours) |
@@ 3 Adaptive Normalization-Free Feature Recalibration
3.1 Background and Notation
We consider a FL setting with clients, each owning a dataset of image-label pairs and optimizing a local objective , where is a loss function and the model parameters. Heterogeneity among can degrade the global model performance and slow convergence (Kairouz et al., 2021). In this study, we modify the backbone model to address this. As they are the most widely used family, and they perform better or on par with others (Pieri et al., 2023; Siomos et al., 2024), we focus specifically on convolutional neural networks (CNNs). Let represent a batch of image samples with channels and dimensions . For a convolutional layer with weights and a kernel size of , the outputs are given by:
with the dimensions of being and those of . In typical CNNs, the activations are then normalized:
where are learnable parameters, is an indexing vector and is the set of pixels over which are computed. BN computes statistics along the axes, LN along , and GN along separately for each of groups of channels. Channel Attention (CA) mechanisms, like the Squeeze-and-Excitation (SE) block (Hu et al., 2018), recalibrate feature responses by modeling inter-channel relationships. The channel descriptor is obtained via Global Average Pooling (GAP):
This descriptor is then non-linearly transformed to capture dependencies between channels; in SE blocks this is done via the learnable weights and , where is a dimensionality reduction ratio: , where , is the sigmoid function and the ReLU function, yielding per-channel scaling factors which are applied to the normalized activations .
3.2 Effect of Normalization on Channel Attention
@@ 3.3 Adaptive Normalization-Free Feature Recalibration
Comparing (4) and (6), we note several advantages of ANFR. First, and are computed from convolutional weights, not the activations. Since weights are initialized identically and synchronized during FL, these weight-derived statistics are consistent across clients. Moreover, the second term of (6) now captures statistics of the input before convolution, providing an additional calibration point for CA and bypassing the effect of . By applying CA after SWS, we ensure channel descriptors are not distorted by batch-dependent statistics or cross-channel dependencies introduced by activation normalization. This allows CA to adjust channel responses effectively, improving the model’s capacity to learn stable feature representations that are consistent across clients with diverse data distributions. Therefore, the combination of SWS and CA overcomes the drawbacks of traditional normalization methods in federated learning, providing a novel and effective solution for improving model performance in the presence of data variability. Lastly, we note ANFR operates at the model level and inherits the theoretical convergence guarantees of the aggregation method it is used with.
3.4 Mechanistic Interpretability Analysis
@@
Next, we conduct a mechanistic interpretability analysis comparing the effects of BN and SWS on class selectivity and attention weight variability to further substantiate the effectiveness of integrating CA with SWS. We examine how well the ANFR model discriminates between classes before111All networks are pre-trained on ImageNet. and after training on the heterogeneous ‘split-3’ partitioning of CIFAR-10 from Qu et al. (2022). This evaluation helps understand how our method improves class discriminability under data heterogeneity. We isolate the effect of different components by comparing ANFR (using SWS with CA), BN-ResNet (using BN), NF-ResNet (using SWS without CA), and SE-ResNet (using CA with BN). Class selectivity is quantified by the class selectivity index (CSI) (Morcos et al., 2018), defined for each neuron as , where is the class-conditional activation that elicits the highest response and is the mean activation for all other classes. A right-skewed CSI distribution indicates higher class selectivity, crucial for effective classification under heterogeneous data. Lastly, we examine the distribution of attention weights, like done in Wang et al. (2020), for models using CA, to understand its contribution to class discrimination.
@@
Figure 3 shows CSI distributions for the last layer before the classifier, where class specificity is maximized in CNNs. Before FL training, incorporating CA in SE-ResNet slightly increases class selectivity compared to BN-ResNet. Combining CA with SWS in ANFR shows negligible change in class selectivity compared to NF-ResNet, indicating CA’s minimal impact at this stage. However, after training on heterogeneous data, we observe a notable shift: BN reduces class selectivity (compared to before training), evidenced by left-skewed distributions for BN-ResNet and SE-ResNet. Adding CA increases class selectivity for both normalization methods, but due to receiving inconsistently normalized inputs (4) cannot fully mitigate BN’s negative effect. The ANFR model, however, shows a significant increase in class selectivity compared to NF-ResNet, with strong class selectivity () units nearly doubling from to . This improvement manifests only after FL training, indicating that combining CA and SWS in ANFR enhances the model’s ability to specialize and discriminate classes under data heterogeneity.
@@
In Figure 3 we use the variability of attention weights across channels and classes as an indicator of adaptation: high variability suggests CA is actively re-weighing features to adapt to different class characteristics. Before FL training (left panel), both SE and ANFR models display high variability, as, when heterogeneity is not a factor, CA provides a diverse and informative signal for both activation and weight normalization. After FL training (right panel), the attention mechanism of SE-ResNet turns into an identity operator, with attention weights converging to across all channels and classes, meaning SE-ResNet fails to preserve the discriminative power of CA under heterogeneity. In contrast, ANFR maintains high variability in CA weights across channels and classes. This sustained variability implies that CA remains active and continues to provide class-discriminative signals when combined with weight standardization.
@@
These insights support our design choices. BN’s adverse effects in heterogeneous FL are highlighted by diminished class selectivity and inactive CA in SE-ResNet, while ANFR maintains and improves class selectivity, demonstrating that integrating CA with weight standardization effectively counters data heterogeneity. The enhanced class selectivity in ANFR correlates with improved downstream performance in heterogeneous FL settings, as we show in Section 4. Additional details and extended CSI and attention weight results from other layers are presented in Appendix D.